Ammonia is produced by the reaction of nitrogen and hydrogen.
N2(g)+3H2(g)→2NH3(g)Ammonia

How many moles of NH3 are produced when 1.2 mol of H2 reacts?
Express your answer to two significant figures and include the appropriate units.
1
Expert's answer
2015-10-20T04:00:33-0400
Solution
N(NH3) = 2/3 n(H2) = 2/3 * 1.2 = 0.8 (mol)

Answer
N(NH3) = 0.80 mol